A car travels along a horizontal circular curved
road that has a radius of 600 m. If the speed is uniformly
increased at a rate of 2000 km>h2, determine the magnitude
of the acceleration at the instant the speed of the car is
60 km>h.

Grade:Upto college level

2 Answers

Yogita Bang
11 years ago

a2 = (v2/r)2 + aT = (3600/0.6)2 + (2000)2 = (6000)2 + (2000)2 = 40*106

a≈6.32*103 km/h2

lokesh soni
11 years ago

a2=(ar)2+(at)2

a2=v2/r+(2000km/hr2)2

a2=36000000+4000000

a=6.32*103